# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ 5.1.4 Inference
$\frac{\widehat{\theta}_{1}-\theta_{1}^{\star}}{\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1\right).$
For simplicity we set both $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ to be all zero vectors. In Figure 7, we present the QQ-plot comparing the generated instances and the theoretical distributions. In both cases, the spread of the term of interest matches the normality claimed in Theorem 3 and Corollary 1.
Figure 7: Quantile-quantile plots of $\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}$ and $(\widehat{\theta}_{1}-\theta_{1}^{\star})/\sqrt{[\bm{L}_{Lz}^{\dagger}]_{11}}$ v.s. zero-mean normal distributions. The estimate $\widehat{\bm{\theta}}$ is computed with $\mathsf{RP\text{-}MLE}$. The parameters are set to be $n=10000,m=10,p=0.5$. Data is collected over 10000 trials.
Coverage of the confidence intervals.
We compare the empirical coverage rate of the two-sided confidence intervals in Proposition 3 with the theoretical ones. Without loss of generality we consider the confidence interval for the first coordinate. In each trial, we compute the empirical coverage rate
$1-\widehat{\alpha}=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\{\mathcal{C}_{i}^{-}(\alpha/2)\leq\theta_{i}^{\star}\leq\mathcal{C}_{i}^{+}(\alpha/2)\}$
and report the average $1-\widehat{\alpha}$ over the trials. We do this for multiple $\alpha$’s and compare $1-\alpha$ with $1-\widehat{\alpha}$ in Table 1. The empirical coverage rate $1-\widehat{\alpha}$ is very close to the theoretical coverage rate $1-\alpha$.
$1-\alpha$ 0.8 0.9 0.95 0.98 0.99 0.995 0.999
$1-\widehat{\alpha}$ 0.801145 0.900635 0.95108 0.980395 0.99025 0.99523 0.99904
Table 1: Theoretical coverage rate $1-\alpha$ and empirical coverage rate $1-\widehat{\alpha}$ of the two-sided confidence intervals $[\mathcal{C}_{i}^{-}(\alpha/2),\mathcal{C}_{i}^{+}(\alpha/2)]$. The parameters are set to be $n=10000,m=20,p=0.1$. The result is averaged over $10000$ trials.
5.2 LSAT real data
We study a real dataset (LSAT) on the Law School Admissions Test from [DBL70]. LSAT has full observation of 1000 people answering 5 problems. Foro each person-item pair it records whether the answer is correct or not. The second row in Table 2 lists how many people answer each problem correctly. From the first look, Problem 3 seems to be the hardest question. We proceed to quantify the hardness of these problems under the Rasch model and infer how confident we are in claiming it is the hardest. We use $\mathsf{RP\text{-}MLE}$ to get an estimate of the latent scores and calculate two-sided confidence intervals at level $\alpha=0.01$ for each coordinate with the framework in Section 3.3. See Table 2 for the result. Higher latent score here means higher difficulty. The estimated parameters agree with the total number of correct answers in reverse order. We can also see that the confidence interval lower bound of $\theta_{3}^{\star}$ is larger than the largest confidence interval upper bound of $\theta_{1}^{\star},\theta_{2}^{\star},\theta_{4}^{\star},\theta_{5}^{\star}$. With Bonferroni correction, we have a 0.95 confidence level to say Problem 3 is the most difficult problem.
Problem 1 2 3 4 5
Total correct 924 709 553 763 870
$\theta$ estimate -1.2139 0.4093 1.1123 0.2977 -0.6053
CI lower bound -1.5957 0.1449 0.8321 0.0234 -0.9050
CI upper bound -0.8322 0.6736 1.3926 0.5720 -0.3057
Table 2: Latent score estimate calculated using $\mathsf{RP\text{-}MLE}$ and confidence interval calculated with the framework in Section 3.3. The significance level is chosen to be $\alpha=0.01$ for each coordinate.
Now we assume Problem 3 is the top-$1$ item and investigate the top-$1$ recovery rate of different algorithms on LSAT under incomplete observation. In each trial we randomly select $\widetilde{n}$ people, and for each of them we randomly select their outcome on $\widetilde{m}$ problems. We then estimate $\bm{\theta}^{\star}$ using the subsampled data with different methods and compare their top-1 recovery rate, i.e., the proportion of trials where the top-$1$ item is correctly identified. In Figure 8 we see results similar to the simulation. Our algorithm $\mathsf{MRP\text{-}MLE}$ has a similar recovery rate compared to PMLE and spectral method.
Figure 8: Top-1 recovery rate using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameters are chosen to be $\widetilde{m}=4$ and $\widetilde{n}$ varies from 100 to 200. The result is averaged over 10000 trials.
6 Discussion
In this paper, we propose two new likelihood-based estimators $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ for item parameter estimation in the Rasch model. Both enjoy optimal finite sample estimation guarantee and $\mathsf{RP\text{-}MLE}$ is further amenable to tight uncertainty quantification. All this is achieved even when the user-item response data are extremely sparse (cf. [NZ23]). Below, we identify several questions that are interesting for further investigation:
  • Does PMLE achieve optimal theoretical guarantee? In our experiments, pseudo MLE has shown a similar performance to $\mathsf{MRP\text{-}MLE}$. This naturally leads to the question of whether PMLE can enjoy the same theoretical guarantee. This is relevant to our work because our methods can be viewed as a modification of pseudo MLE by incorporating random disjoint pairing to decouple statistical dependency among paired $Y_{ij}$’s. It remains unclear whether such dependency is a fundamental bottleneck.
  • Inference for multiple data splitting. In $\mathsf{MRP\text{-}MLE}$ (Algorithm 2), we average the estimates over multiple runs of data splitting to improve the empirical result over $\mathsf{RP\text{-}MLE}$. While this is effective for estimation, we do not have an immediate analog for inference as the estimations $\{\widehat{\bm{\theta}}^{(i)}\}_{i\in[n_{\mathrm{split}}]}$ are dependent on each other. It is an interesting question whether we can characterize the distribution of $\widehat{\bm{\theta}}$ in $\mathsf{MRP\text{-}MLE}$ and whether it leads to better inference, e.g., tighter confidence intervals.
  • Extending random pairing to other models in IRT. Some IRT models parameterize the latent score of users and items differently from the Rasch model. For instance, the two-parameter logistic model (2PL) assumes that $X_{ti}$, the response of user $t$ to item $i$, follows the law
    $\mathbb{P}[X_{ti}=1]=\frac{\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})}{1+\exp(a_{t}^{\star}+b_{t}^{\star}\theta_{i}^{\star})},$
    where $\bm{\theta}^{\star}$ is the latent score of the items while $\bm{a}^{\star}$ and $\bm{b}^{\star}$ are two parameter vectors for the users. Unlike the Rasch model, in the 2PL model,
    $\mathbb{P}[X_{ti}>X_{tj}\mid X_{ti}\neq X_{tj}]=\frac{\exp(b_{t}^{\star}\theta_{i}^{\star})}{\exp(b_{t}^{\star}\theta_{i}^{\star})+\exp(b_{t}^{\star}\theta_{j}^{\star})}$
    is not independent of the user parameter $\bm{b}^{\star}$. Therefore reduction to the BTL model is no longer true in this case. It is interesting and non-trivial to extend the idea of random pairing to the 2PL model.
Acknowledgement
YY and CM were partially supported by the National Science Foundation via grant DMS-2311127.
Appendix A Proof of Fact 1
Expanding the probability of the random events, we have that
$\mathbb{P}[X_{ti}<X_{tj}\mid X_{ti}\neq X_{tj}]$ $=\frac{\mathbb{P}[X_{ti}=0,X_{tj}=1]}{\mathbb{P}[X_{ti}=0,X_{tj}=1\text{\text{ or }}X_{ti}=1,X_{tj}=0]}$
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\Bigg{/}\left(\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}+\frac{e^{\theta_{i}^{\star}}e^{\zeta_{t}^{\star}}}{(e^{\zeta_{t}^{\star}}+e^{\theta_{i}^{\star}})(e^{\zeta_{t}^{\star}}+e^{\theta_{j}^{\star}})}\right)$
$=\frac{e^{\zeta_{t}^{\star}}e^{\theta_{j}^{\star}}}{e^{\zeta_{t}^{\star}}(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})}$
$=\frac{e^{\theta_{j}^{\star}}}{e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}}}.$
Now consider $\mathbb{P}[X_{ti}\neq X_{tj}]$, we have
Let $f:[1/\kappa_{2},\kappa_{2}]^{2}\rightarrow\mathbb{R}$ defined by
$f(a,b)\coloneqq\frac{a+b}{(1+a)(1+b)}.$
Its partial derivatives are
@@ B.1 Degree range of $\mathcal{G}_{X}$
Recall that $n_{t}$ is the number of neighbors of user $t$ in $\mathcal{G}_{X}$. Furthermore, we denote $m_{i}$ as the number of users that is compared with problem $i$ and at least another item, i.e.,
$m_{i}\coloneqq\left|\{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\}\right|.$ (17)
The following lemma controls the size of $n_{t}$ and $m_{i}$.
@@
$n_{t}\leq\left(\frac{3}{2}mp\right)\vee 165\log(n).$ (19)
@@
$\mathbb{P}\left[t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2\right]$ $=\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]-\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\text{ and }n_{t}=1\right]$
@@
as long as $mp\geq 2$. Let $\mu_{i}\coloneqq\mathbb{E}[m_{i}]$. By the linearity of expectation, we have
@@
$\mathbb{P}[|m_{i}-\mu_{i}|\leq(1/2)\mu_{i}]$ $\leq 2e^{-\frac{1}{12}\mu_{i}}\leq 2e^{-\frac{1}{24}np}\leq m^{-1}n^{-10}$
@@
$\mathbb{P}[n_{t}\geq(3/2)mp]\leq e^{-\frac{1}{10}mp}\leq n^{-11}.$
In the case of $mp<110\log(n)$, the quantity $\mathbb{P}\left[n_{t}\geq 165\log(n)\right]$ clear decreases as $p$ decreases. So we may use the case $mp=110\log(n)$ to bound this quantity and conclude that
$\mathbb{P}\left[n_{t}\geq 165\log(n)\right]\leq n^{-11}.$
@@ B.2 Proof of Lemma 2
=$\overset{(\text{i})}{=}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}L_{ij}^{t}$
$\overset{(\text{ii})}{\leq}\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\sum_{j:j\neq i}R_{ij}^{t}\leq m_{i}\leq\frac{3}{2}np.$
Here (i) holds since $L_{ij}^{t}$ can only be 0 when $n_{t}\leq 1$, and (ii) holds since $L_{ij}^{t}\leq R_{ij}^{t}$ by definition. For the lower bound of $d_{i}$, notice that for any $(t,i)$, $\sum_{j}L_{ij}^{t}$ is either 00 or 1. Fix $\mathcal{E}_{X}$ and only consider randomness on $L_{ij}^{t}$. By Hoeffding’s inequality,
$\mathbb{P}\left\{d_{i}-\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j:j\neq i}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\leq-\frac{1}{12\kappa_{2}}np\right\}$ $\leq\exp\left(-\frac{(1/72)\kappa_{2}^{-2}n^{2}p^{2}}{m_{i}}\right)$
@@
$\geq\frac{2\lfloor n_{t}/2\rfloor}{n_{t}}\cdot\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq\frac{1}{3\kappa_{2}}$
as long as $n_{t}\geq 2$. The first inequality here uses Fact 1. Then by definition of $m_{i}$ in (17),
$\sum_{t:(t,i)\in\mathcal{E}_{X},n_{t}\geq 2}\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\geq\frac{1}{3\kappa_{2}}m_{i}.$ (23)
@@
$d_{i}\geq\frac{1}{3\kappa_{2}}m_{i}-\frac{1}{12\kappa_{2}}np\geq\frac{1}{24\kappa_{2}}np.$
@@ B.3 Proof of Lemma 3
For the upper bound, it is clear from Lemmas 10 and 2 that $\lambda_{1}(\bm{L}_{L})\leq 2\max_{i}d_{i}\leq 3np$. For the lower bound, we use the matrix Chernoff inequality (see Section 5 of [Tro15]). Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then $\lambda_{m-1}(\bm{L}_{L})=\lambda_{m-1}(\bm{R}\bm{L}_{L}\bm{R}^{\top})$. For any $t\in[n]$, by (19),
$0\leq\lambda_{m-1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})\leq\lambda_{1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})=\lambda_{1}(\bm{L}_{L}^{t})\leq 2$
@@
Moreover $\sum_{i>j}\mathbb{E}R_{ij}^{t}\geq\frac{1}{2}(\mathbb{E}[n_{t}]-1)=(mp-1)/2$, where the $-1$ accounts for possible unpaired $X_{ti}$. By symmetry $\mathbb{E}R_{ij}^{t}$ is the same for any $(i,j)$. Then for any $(i,j)$,
@@
Appendix C Proof of Lemma 4
@@
We study the MLE $\widehat{\bm{\theta}}$ by analyzing the iterates of preconditioned gradient descent starting from the ground truth. Let $\bm{\theta}_{0}=\bm{\theta}^{\star}$ be the starting point and $\eta>0$ be the stepsize that is small enough. The preconditioned gradient descent iterates are given by
$\bm{\theta}^{t+1}=\bm{\theta}^{t}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{t}).$ (24)
Consider the Taylor expansion of $\nabla\mathcal{L}(\bm{\theta}^{t})$, we have
$\nabla\mathcal{L}(\bm{\theta}^{t})$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-Y_{ji}^{(l)}+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})+\sigma(\theta_{i}^{t}-\theta_{j}^{t})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
@@
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\left(\left(-\epsilon_{ji}^{(l)}+\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})(\delta_{i}^{t}-\delta_{j}^{t})+\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}\right)(\bm{e}_{i}-\bm{e}_{j})\right).$
@@
Here $\bm{\delta}^{t}\coloneqq\bm{\theta}^{t}-\bm{\theta}^{\star}$ and for all $(i,j)$, $\xi_{ij}^{t}\in\mathbb{R}$ is some number that lies between $\theta_{i}^{\star}-\theta_{j}^{\star}$ and $\theta_{i}^{t}-\theta_{j}^{t}$. As $\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=z_{ij}$ and $\delta_{i}^{t}-\delta_{j}^{t}=(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{\delta}^{t}$, we have
$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}\bm{\delta}^{t}-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
@@
where $\bm{r}^{t}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}[\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{t})(\delta_{i}^{t}-\delta_{j}^{t})^{2}(\bm{e}_{i}-\bm{e}_{j})]$. Feeding this into (24), we have
$\bm{\delta}^{t+1}=\left(1-\eta\right)\bm{\delta}^{t}-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\right)$ (25)
@@
By definition $\bm{\delta}^{0}=\bm{0}$. Applying this recursive relation $t-1$ times we obtain
$\bm{\delta}^{t}$ $=-\eta\sum_{i=0}^{t-1}(1-\eta)^{i}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}$
$=-\left[1-(1-\eta)^{t}\right]\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{t-1}(1-\eta)^{t-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}.$
@@
At this point, we invoke an existing result on these terms that have been studied in [YCOM24] for a more general setting. The proof of Lemma 2 combined with Lemmas 3 and 4 in [YCOM24] reveal the following properties of (24).

### v2

@@ A.4 Analysis for asymptotic normality
$=\left[\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\right](\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}).$
Here the second row comes from the optimality condition of $\widehat{\bm{\theta}}_{k}^{(n)}$. Now under some regularity conditions, we can use the consistency of $\widehat{\bm{\theta}}_{k}^{(n)}$ to show
$\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\mathrm{d}\tau\approx\bm{H}^{\infty}.$ (20)
Then $\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)}\approx(\bm{H}^{\infty})^{\dagger}\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$. Note that $\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ is zero-mean and independent between different user $t$. This independence also holds for $\sum_{k=1}^{n_{\mathrm{split}}}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ and $\nabla\mathcal{L}_{\mathrm{WP}}^{(t)}(\bm{\theta}^{\star})$. Then we can invoke central limit theorem to reach the desired result.
@@ B.1 Degree range of $\mathcal{G}_{X}$
Recall that $m_{t}$ is the number of neighbors of user $t$ in $\mathcal{G}_{X}$. Furthermore, we denote $n_{i}$ as the number of users that is compared with problem $i$ and at least another item, i.e.,
$n_{i}\coloneqq\left|\{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2\}\right|.$ (21)
The following lemma controls the size of $m_{t}$ and $n_{i}$.
@@
$m_{t}\leq\left(\frac{3}{2}mp\right)\vee 165\log(n).$ (23)
@@
$\mathbb{P}\left[t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2\right]$ $=\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\right]-\mathbb{P}\left[(t,i)\in\mathcal{E}_{X}\text{ and }m_{t}=1\right]$
@@
as long as $mp\geq 2$. Let $\mu_{i}\coloneqq\mathbb{E}[n_{i}]$. By the linearity of expectation, we have
@@
$\mathbb{P}[|n_{i}-\mu_{i}|\leq(1/2)\mu_{i}]$ $\leq 2e^{-\frac{1}{12}\mu_{i}}\leq 2e^{-\frac{1}{24}np}\leq m^{-1}n^{-10}$
@@
$\mathbb{P}[m_{t}\geq(3/2)mp]\leq e^{-\frac{1}{10}mp}\leq n^{-11}.$
In the case of $mp<110\log(n)$, the quantity $\mathbb{P}\left[m_{t}\geq 165\log(n)\right]$ clear decreases as $p$ decreases. So we may use the case $mp=110\log(n)$ to bound this quantity and conclude that
$\mathbb{P}\left[m_{t}\geq 165\log(n)\right]\leq n^{-11}.$
@@ B.2 Proof of Lemma 2
$\overset{(\text{i})}{=}\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\sum_{j:j\neq i}L_{ij}^{t}$
$\overset{(\text{ii})}{\leq}\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\sum_{j:j\neq i}R_{ij}^{t}\leq n_{i}\leq\frac{3}{2}np.$
Here (i) holds since $L_{ij}^{t}$ can only be 0 when $m_{t}\leq 1$, and (ii) holds since $L_{ij}^{t}\leq R_{ij}^{t}$ by definition. For the lower bound of $d_{i}$, notice that for any $(t,i)$, $\sum_{j}L_{ij}^{t}$ is either 0 or 1. Fix $\mathcal{E}_{X}$ and only consider randomness on $L_{ij}^{t}$. By Hoeffding’s inequality,
$\mathbb{P}\left\{d_{i}-\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\mathbb{E}\left[\sum_{j:j\neq i}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\leq-\frac{1}{12\kappa_{2}}np\right\}$ $\leq\exp\left(-\frac{(1/72)\kappa_{2}^{-2}n^{2}p^{2}}{n_{i}}\right)$
@@
$\geq\frac{2\lfloor m_{t}/2\rfloor}{m_{t}}\cdot\frac{2\kappa_{2}}{(1+\kappa_{2})^{2}}\geq\frac{1}{3\kappa_{2}}$
as long as $m_{t}\geq 2$. The first inequality here uses Fact 1. Then by definition of $n_{i}$ in (21),
$\sum_{t:(t,i)\in\mathcal{E}_{X},m_{t}\geq 2}\mathbb{E}\left[\sum_{j}L_{ij}^{t}\mid(t,i)\in\mathcal{E}_{X}\right]\geq\frac{1}{3\kappa_{2}}n_{i}.$ (27)
@@
$d_{i}\geq\frac{1}{3\kappa_{2}}n_{i}-\frac{1}{12\kappa_{2}}np\geq\frac{1}{24\kappa_{2}}np.$
@@ B.3 Proof of Lemma 3
For the upper bound, it is clear from Lemmas 12 and 2 that $\lambda_{1}(\bm{L}_{L})\leq 2\max_{i}d_{i}\leq 3np$. For the lower bound, we use the matrix Chernoff inequality (see Section 5 of [43]). Let $\bm{R}\in\mathbb{R}^{(n-1)\times n}$ be a partial isometry such that $\bm{R}\bm{R}^{\top}=\bm{I}_{m-1}$ and $\bm{R}\bm{1}=\bm{0}$. Then $\lambda_{m-1}(\bm{L}_{L})=\lambda_{m-1}(\bm{R}\bm{L}_{L}\bm{R}^{\top})$. For any $t\in[n]$, by (23),
$0\leq\lambda_{m-1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})\leq\lambda_{1}(\bm{R}\bm{L}_{L}^{t}\bm{R}^{\top})=\lambda_{1}(\bm{L}_{L}^{t})\leq 2.$
@@
Moreover $\sum_{i>j}\mathbb{E}R_{ij}^{t}\geq\frac{1}{2}(\mathbb{E}[m_{t}]-1)=(mp-1)/2$, where the 1-1 accounts for possible unpaired $X_{ti}$. By symmetry $\mathbb{E}R_{ij}^{t}$ is the same for any $(i,j)$. Then for any $(i,j)$,
@@
Appendix C Proofs for Section 3.2
In this section, we provide the proofs for the results related to the non-asymptotic expansion in Section 3.2.
C.1 Proof of Theorem 2
@@
We study the MLE $\widehat{\bm{\theta}}$ by analyzing the iterates of preconditioned gradient descent starting from the ground truth. Let $\bm{\theta}_{0}=\bm{\theta}^{\star}$ be the starting point and $\eta>0$ be the stepsize that is small enough. At iteration $\tau$, the preconditioned gradient descent update is given by
$\bm{\theta}^{\tau+1}=\bm{\theta}^{\tau}-\eta\bm{L}_{Lz}^{\dagger}\nabla\mathcal{L}(\bm{\theta}^{\tau}).$ (28)
Recall the definitions: $\sigma(x)=e^{x}/(1+e^{x})$ is the sigmoid function. We also have $z_{ij}\coloneqq e^{\theta_{i}^{\star}}e^{\theta_{j}^{\star}}/(e^{\theta_{i}^{\star}}+e^{\theta_{j}^{\star}})^{2}$, $\widehat{z}_{ij}\coloneqq e^{\hat{\theta}_{i}}e^{\hat{\theta}_{j}}/(e^{\hat{\theta}_{i}}+e^{\hat{\theta}_{j}})^{2}$ and $\epsilon_{ij}^{t}\coloneqq Y_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$. The total number of observed effective comparisons in $\mathcal{G}_{Y}$ is $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$.
We have defined $\bm{\bm{B}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ and $\widehat{\bm{\epsilon}}\in\mathbb{R}^{L_{\mathrm{total}}}$ as
$\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\text{$\quad$(repeat $L_{ij}$ times for edge $(i,j)$)}$
and
$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{t}/\sqrt{z_{ij}},\cdots\right]_{(i,j,t):i>j,(i,j)\in\mathcal{E}_{Y},L_{ij}^{t}=1}\in\mathbb{R}^{L_{\mathrm{total}}}.$
Moreover, define a weighted graph Laplacian
$\bm{L}_{L\tilde{z}}\coloneqq\sum_{(i,j)\in\mathcal{E}_{Y},i>j}L_{ij}\widetilde{z}_{ij}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}$
for $\tilde{z}$ being $z$ or $\widehat{z}$, and $\bm{L}_{L\tilde{z}}^{\dagger}$ is its pseudo-inverse.
Consider the Taylor expansion of $\nabla\mathcal{L}(\bm{\theta}^{\tau})$, we have
$\nabla\mathcal{L}(\bm{\theta}^{\tau})$ $=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-Y_{ji}^{t}+\sigma(\theta_{i}^{\tau}-\theta_{j}^{\tau})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-\epsilon_{ji}^{t}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})+\sigma(\theta_{i}^{\tau}-\theta_{j}^{\tau})\right)(\bm{e}_{i}-\bm{e}_{j})\right)$
@@
$=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{t:L_{ij}^{t}=1}\left(\left(-\epsilon_{ji}^{t}+\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})(\delta_{i}^{\tau}-\delta_{j}^{\tau})+\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{\tau})(\delta_{i}^{\tau}-\delta_{j}^{\tau})^{2}\right)(\bm{e}_{i}-\bm{e}_{j})\right).$
@@
Here $\bm{\delta}^{\tau}\coloneqq\bm{\theta}^{\tau}-\bm{\theta}^{\star}$ and for all $(i,j)$, $\xi_{ij}^{\tau}\in\mathbb{R}$ is some number that lies between $\theta_{i}^{\star}-\theta_{j}^{\star}$ and $\theta_{i}^{\tau}-\theta_{j}^{\tau}$. As $\sigma^{\prime}(\theta_{i}^{\star}-\theta_{j}^{\star})=z_{ij}$ and $\delta_{i}^{\tau}-\delta_{j}^{\tau}=(\bm{e}_{i}-\bm{e}_{j})^{\top}\bm{\delta}^{\tau}$, we can rewrite $\nabla\mathcal{L}(\bm{\theta}^{\tau})$ as
$\nabla\mathcal{L}(\bm{\theta}^{\tau})=\bm{L}_{Lz}\bm{\delta}^{\tau}-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{\tau},$
@@
where $\bm{r}^{\tau}=\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}\cdot[\frac{1}{2}\sigma^{\prime\prime}(\xi_{ij}^{\tau})(\delta_{i}^{\tau}-\delta_{j}^{\tau})^{2}(\bm{e}_{i}-\bm{e}_{j})]$. Feeding this into (28), we have
$\bm{\delta}^{\tau+1}=\left(1-\eta\right)\bm{\delta}^{\tau}-\eta\left(\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}-\bm{L}_{Lz}^{\dagger}\bm{r}^{\tau}\right)$ (29)
@@
By definition $\bm{\delta}^{0}=\bm{0}$. Applying this recursive relation $\tau-1$ times we obtain
$\bm{\delta}^{\tau}$ $=-\eta\sum_{i=0}^{\tau-1}(1-\eta)^{i}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}$
$=-\left[1-(1-\eta)^{\tau}\right]\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\sum_{i=0}^{\tau-1}(1-\eta)^{\tau-1-i}\bm{L}_{Lz}^{\dagger}\bm{r}^{i}.$
@@
At this point, we invoke an existing result on these terms that have been studied in [47] for a more general setting. The proof of Lemma 6 combined with Lemmas 7 and 8 in [47] reveal the following properties of (28). The proof is deferred to Section C.1.1.